# Identity-Obscured Neural Radiance Fields: Privacy-Preserving 3D Facial Reconstruction

### v1

@@ 7 Implementations
In the experimental section, for each group of data samples in the FaceScape dataset, we have 30 original RGB images taken from different camera perspectives, with the shooting perspective being consistent within each group. These images consist of two sets: one comprises privacy-neutral images directly usable for reconstruction, while the other contains sensitive information requiring privacy protection. In Figure 14, we showcase a complete set of image data used for the experiments, which intuitively represents our classification criteria.
@@
Figure 14: We have proposed a reconstruction method based on identity-obscured input. For privacy-neutral images that inherently do not contain sensitive information (Row 1 and Row 2), we retain the RGB images. For other images, we undergo privacy-preserving processing before proceeding with subsequent operations.
@@
Moving on to the second stage, we trained for an additional 500 epochs specifically focusing on reconstructing facial details. This phase involved using identity-obscured inputs and corresponding supervision to achieve a fine-tuning of the geometry, ensuring a more detailed representation of facial features.
Loss Setups.
@@
(a) VolSDF (b) Ours (c) VolSDF (d) GT
Template Acquisition.
We trained a set of head templates using an extensive dataset, with each template constructed from 10 unique identity heads. These templates were obtained through the method proposed by Xu et al. [37]. The training process resulted in some template heads geometry, as exemplified in Figure 16. Importantly, these templates are neutral and do not contain any sensitive information. The representation of a template facilitates its direct integration into the training process of stage 2, where privacy-protected inputs are utilized.
(a) (b)
Figure 16: Template: neutral geometric representation of a head devoid of sensitive information, usable to support stage 2 training.
@@
Resolution Impact.
It is important to note that, for the sake of privacy protection, we actively engage in supervised reconstruction of all our images at a 64 resolution level. Despite the impact of resolution reduction on reconstruction quality, this effect remains unrelated to the robustness of our methodology. In Figure 17, we depict the reconstruction outcomes obtained at both high resolution (512) and low resolution (64) through the utilization of identity-obscured input. This confirmation solidifies the understanding that the observed decline in geometric accuracy, which does not affect identity recognition, is due to the supervision based on low-resolution images.
8 Future Work.
The introduction of NeRF has elevated the precision of geometric reconstruction, and substantial efforts have been dedicated to minimizing the reconstruction cost and time associated with NeRF [20, 3, 14]. We believe that the task of facial reconstruction using NeRF may have broader applications. For instance, numerous recognition systems could potentially discard the supervision of original 2D features and adopt the truly 3D-reconstructed facial geometry as the ID identifier.
Our approach introduces a potential avenue for privacy protection. In fact, as the importance of privacy preservation continues to grow, it is likely that more privacy protection methods will be proposed and integrated into our framework. By avoiding direct exposure of privacy information during usage, we have demonstrated the feasibility of achieving high-fidelity geometric reconstruction through identity-obscured input. This finding inspires confidence that future endeavors can advance both security and ethics through new explorations in this direction.

### v2

@@ Conclusion
In the experimental section, for each group of data samples in the FaceScape dataset, we have 30 original RGB images taken from different camera perspectives, with the shooting perspective being consistent within each group. These images consist of two sets: one comprises neutral images directly usable for reconstruction, while the other contains sensitive information requiring protection (see Figure 17). In Figure 18, we showcase a complete set of image data used for the experiments from the FaceScape dataset, which intuitively represents our classification criteria.
Figure 17: An example of facial image sensitivity classification from the FaceScape dataset.
@@
Figure 18: We have proposed a reconstruction method based on non-sensitive input. For neutral images that inherently do not contain sensitive information (Row 1,2), we retain the RGB images. For other images, we undergo protected processing before proceeding with subsequent operations(Row 3,4,5,6).
(a) VolSDF (b) Ours (c) VolSDF (d) GT
@@
Moving on to the second stage, we trained for an additional 500 epochs specifically focusing on reconstructing facial details. This phase involved using non-sensitive inputs and corresponding supervision to achieve a fine-tuning of the geometry, ensuring a more detailed representation of facial structures.
Loss Setups
@@
Template Acquisition
We trained a set of head templates using an extensive dataset, with each template constructed from 10 unique identity heads. These templates were obtained through the method proposed by Xu et al. (Xu et al. 2023). The training process resulted in some template head geometry. Importantly, these templates are neutral and do not contain any sensitive information. The representation of a template facilitates its direct integration into the training process of the second stage, where protected inputs are utilized.
Resolution Impact
@@
It is important to note that, for the sake of sensitivity protection, we actively engage in the supervised reconstruction of all our images at a 64-resolution level. Despite the impact of resolution reduction on reconstruction quality, this effect remains unrelated to the robustness of our methodology. In Figure 20, we depict the reconstruction outcomes obtained at both high resolution (512) and low resolution (64) through the utilization of non-sensitive input. This confirmation solidifies the understanding that the observed decline in geometric accuracy, which is within acceptable limits, is due to the supervision based on low-resolution images. To ensure fair experiments, we maintain the same training epochs (1.5k) and employ $512\times 512$ resolution matching cubes to extract the final mesh. The mesh evaluation is conducted by appropriately cropping it to the size of the provided GT mesh, ensuring a fair assessment of facial geometric reconstruction accuracy.
Acknowledgments
This work was supported in part by the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22) and the RIE2020 Industry Alignment Fund–Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).